# Un-mixing Test-time Adaptation under Heterogeneous Data Streams

### v2

@@ Abstract
Deploying deep models in real-world scenarios remains challenging due to significant performance drops under distribution shifts between training and deployment environments. Test-Time Adaptation (TTA) has recently emerged as a promising solution, enabling on-the-fly model adaptation without access to source data. However, its effectiveness degrades significantly in the presence of complex, mixed distribution shifts – common in practical settings – where multiple latent domains coexist. Adapting under such intrinsic heterogeneity, especially in unlabeled and online conditions, remains an open and underexplored challenge. In this paper, we study TTA under mixed distribution shifts and move beyond conventional homogeneous adaptation paradigms. By revisiting TTA from a frequency-domain perspective, we observe that distribution heterogeneity often manifests in Fourier space – for instance, high-frequency components tend to carry domain-specific variations. This motivates us to perform domain-aware separation using high-frequency texture cues, making diverse shift patterns more tractable. To this end, we propose FreDA, a novel Frequency-based Decentralized Adaptation framework that decomposes globally heterogeneous data into locally homogeneous components in the frequency domain. It further employs decentralized learning and augmentation strategies to robustly adapt under complex, evolving shifts. Extensive experiments across various environments (corrupted, natural, and medical) demonstrate the superiority of our proposed framework over the state-of-the-arts.
Index Terms:
Test-time Adaptation, Transfer Learning
@@ I Introduction
Despite their success, current TTA models often are limited to ideal testing conditions, typically involving homogeneous test samples with similar types of distribution shifts. In reality, distribution shifts are mixed, overlapping, and even conflicting [11, 12, 13, 14, 15]. For instance, photo management software handles diverse corruptions, such as noise, blur, compression; medical imaging platforms manage inconsistencies from varied acquisition methods; and autonomous driving systems face fluctuating conditions like lighting, weather, and road types.
@@
Such complexity poses serious challenges for existing methods. While recent efforts have extended TTA to continually changing environments [5], they typically assume a uniform target domain at each time step, as illustrated in the continual setting in Figure 1 (a). Some methods periodically reset the model to its source pre-trained state [7, 9] to mitigate the cumulative effect of sequential distribution shifts on adaptation. Others down-weight outlier target samples [6, 10] to suppress the impact of abrupt or rare shifts. Although these strategies aim to improve robustness under dynamic conditions, they fall short when facing the entangled / mixed real-world distributions (see Figure 1 (b)). Their shortcomings stem from a shared limitation: they lack the capacity to disentangle or localize domain variations. As a result, they often overfit to dominant patterns or suffer from catastrophic forgetting, failing to be deployed in realistic, heterogeneous data streams.
@@
Figure 1: (a) Illustration of three TTA scenarios with single, continual, and mixed domain shifts. (b) Classification error rate on CIFAR-10-C across different TTA settings: performance of conventional models drops sharply under mixed domain shifts.
@@
Capturing such intrinsic heterogeneity under unlabeled, online conditions is particularly challenging, and effective adaptation in this setting remains an open problem. To tackle this, we propose shifting from coarse, homogeneous adaptation to a fine-grained strategy that explicitly disentangles heterogeneous shifts. Our approach leverages the frequency information to disentangle domain variations, which naturally separates data features across frequency bands: high-frequency components capture fine details like edges and textures, while low-frequency components represent global structures such as shape and illumination (Section IV). By decomposing inputs accordingly, we better characterize distributional variations and identify diverse shifts. Moreover, the Fourier transform operates directly on raw pixel-level inputs, which shields the method from performance degradation under large domain gaps, and also enhancing adaptability and deployment efficiency in real-world applications.
@@
Building upon this insight, we introduce a framework, termed Frequency-based Decentralized Adaptation (FreDA). FreDA begins by partitioning incoming data in the Fourier domain, where high-frequency components are used to group samples with similar shift characteristics. This transforms globally heterogeneous inputs into locally homogeneous subsets before any adaptation takes place. Based on these partitions, FreDA deploys multiple local models, each assigned to a specific subset. These models adapt independently to their respective data streams while periodically synchronizing through parameter exchange. The aggregated parameters form a shared base model, which is then used to reinitialize the local models, ensuring continual and coordinated adaptation across domains. This approach not only alleviates clustering errors via collaborative knowledge sharing but also allows local models to capture diverse distribution shifts. To further enhance robustness, we introduce a novel Fourier-based augmentation scheme that improves sample quality and strengthens adaptation to shift-specific characteristics.
To summarize, the main contributions of this work are three-fold:
@@
  • We identify a key limitation of most existing TTA methods – their neglect of real-world data heterogeneity – which leads to suboptimal performance when confronted with mixed and diverse distribution shifts.
@@
  • We validate FreDA through extensive experiments on corrupted, natural, and medical benchmarks, demonstrating consistent improvements over state-of-the-art methods across various TTA scenarios.
@@ II-A Transfer Learning, Domain Adaptation, Test-time Training, and Test-time Adaptation
Deep neural networks often experience performance degradation when deployed in environments different from their training settings, due to distribution shifts between source (training) and target (testing) domains. This issue is extensively studied under the umbrella of transfer learning [16, 17], which aims to transfer knowledge across domains. A key branch of transfer learning is domain adaptation (DA) [18, 19, 20, 21], where models trained on a labeled source domain are adapted to a target domain. Depending on the availability of source and target data, DA can be classified into: Supervised DA using labeled data in both source and target domains during training; Unsupervised DA, which relies on labeled source data and unlabeled target data; and Source-free DA, that adapts to the target domain without access to source data, typically due to privacy or transmission constraints.
TABLE I: Comparison of different transfer learning settings.
Topic Source Data Target Labels Online Adaptation Model Agnostic
Supervised Domain Adaptation $\checkmark$ $\checkmark$ $\times$
Unsupervised Domain Adaptation $\checkmark$ $\times$ $\times$
Source-free Domain Adaptation $\times$ $\times$ $\times$
Test-time Training $\times$ $\times$ $\checkmark$ $\times$
Test-time Adaptation $\times$ $\times$ $\checkmark$ $\checkmark$
Note: TTA is the most practical setting for real-world deep model deployment.
While traditional DA methods assume access to the whole target distribution during adaptation, such access is not always possible in real-world deployment. To address this, two prominent test-time adaptation paradigms have been proposed: Test-time Training (TTT) [22, 23, 24] introduces an auxiliary task – often self-supervised – during pre-training. At test time, the model optimizes this auxiliary objective to adapt to the target distribution accordingly; Test-time Adaptation (TTA) [3, 7, 8] poses a more demanding and practical scenario, where the model must adapt on-the-fly to the test stream without any prior modification during training. This setting emphasizes the need for rapid, real-time model updates to effectively capture the continuously incoming data. A brief comparison of these adaptation settings is summarized in TABLE I. Our work falls under the TTA setting, where we propose a generalizable and robust adaptation framework that maintains strong performance across diverse distribution shifts and corruptions – without relying on source data or altering the pre-training process.
@@ II-B1 Mixed Domains
Modern TTA methods designed for continual domain shifts [5, 25, 6, 9, 10] have pushed beyond the single adaptation setting. However, these approaches assume that the target domain is uniform at each time point – a special case of mixed domains where the change occurs across distinct time periods. In reality, data streams often involve genuinely mixed domains with heterogeneous distributions present simultaneously. Existing approaches primarily focus on stabilizing model updates through periodic parameter resets [7, 9] or importance weighting of target samples [6, 9, 10], but these techniques are ill-equipped to disentangle and manage the complex interleaving of domain shifts, leading to degraded performance. In contrast, our work adopts a data-centric perspective by explicitly addressing mixed domains via frequency-space decomposition, which facilitates domain separation and proactive distribution alignment prior to model adaptation. While recent efforts like [7] acknowledge mixed distribution challenges under the broader “Dynamic Wild World” paradigm, their unified treatment of diverse real-world factors does not systematically address data heterogeneity. Conversely, our study directly targets the core issue of mixed domains in TTA, introducing tailored strategies for disentangling and adapting to co-occurring heterogeneous distributions.
Figure 2: (a) Visual heterogeneity across five healthcare centers in Camelyon17 dataset: Example patches showcase diverse domain shifts that characterize mixed target domains. (b) Conventional sample latent features from pretrained models fail to separate different target subdomains, showing significant overlap. (c) High-frequency information enable distinct separation of target subdomains, contrasting with the conventional ones.
@@ II-B2 Dependent Sampling
The second challenge stems from class-level dependencies introduced by temporal data sampling. This topic has garnered significant research attention, with a growing body of work [25, 26, 27, 28, 29, 30] actively addressing it through strategies such as pseudo-label-based rebalancing, or extended observation windows. These approaches aim to correct the temporal imbalance caused by skewed class distributions over time. However, unlike this these studies that focus on class-level imbalance, our work diverges by targeting sample-level heterogeneity, which remains underexplored. Specifically, we address style diversity and distributional variation at test-time – a challenge that persist even under class-balanced conditions – to improve model robustness under mixed-domain scenarios.
@@ III Problem Definition of mixed domain shifts
Methods C10 C100 IN
@@
Figure 3: Experimental Result (error rate) and t-SNE feature visualization using TBN [31] as the pseudo-labeling method under mixed distribution shifts, comparing centralized and decentralized adaptation. Different colors represent different classes. (a) Centralized Adaptation: Global BN parameters are applied to the entire batch. (b) Decentralized Adaptation: Localized BN parameters are tailored for clusters separated based on high-frequency features. C10, C100, IN respectively denotes CIFAR-10-C, CIFAR-100-C and ImageNet-C datasets.
IV TTA under Mixed Distribution Shifts: A Fourier Perspective
While Test-Time Adaptation (TTA) methods excel under single-type distribution shifts, their performance degrades catastrophically when facing mixed distribution shifts – a ubiquitous challenge in real-world deployment. As demonstrated on CIFAR-10-C (Figure 1 (b)), leading methods like TENT [3], SAR [7] and RoTTA [25] suffer 16% average accuracy drop compared to single-shift scenarios. Conventional TTA paradigms attempt to mitigate this through passive robustness measures, such as entropy regularization [3], gradient masking [7], or model resetting [9]. These methods treat mixed distribution shifts as irreducible noise, deferring corrective actions until model failures occur – a reactive paradigm that inevitably compromises adaptation efficacy.
To address this, we propose a paradigm shift to proactive distribution management – instead of demanding models to handle arbitrary mixtures, we first resolve the heterogeneity via frequency-space decomposition. Our key insight is that mixed distributions exhibit spectral signatures that can be disentangled before model adaptation begins. As evidenced in Figure 2 (c), high-frequency components naturally separate distinct shifts without relying on error-prone model features (see Figure 2 (b). This theoretically grounded decomposition converts the ill-posed mixed-shift adaptation into multiple well-conditioned homogeneous-shift problems, effectively create an ideal adaptation environment from chaotic real-world streams. Crucially, this decomposition enables a decentralized adaptation where we can deploy domain-specific models tailored to high-frequency-separated clusters. Take TBN [31] as an example: while its centralized variant recalculates batch normalization (BN) statistics globally (treating heterogeneous data as a single distribution), our decentralized adaptation computes localized BN parameters within each spectrally disentangled cluster. As shown in Figu 3, this context-aware estimation preserves semantic discriminability: samples exhibit tighter within-class cohesion and sharper between-class boundaries compared to the blurred separability of centralized adaptation. Building on these insights, the following section details how leveraging the frequency domain enhances TTA methods for realistic scenarios with mixed distribution shifts.
Figure 4: Illustration of Frequency-based Decentralized Learning.
V Frequency-based Decentralized Adaptation
The previous discussion highlights how heterogeneity within target distributions can hinder effective model adaptation. This naturally raises the question: How can we manage such distributional heterogeneity to enable better adaptation? As outlined earlier, distinguishing samples associated with different distribution shifts offers a promising pathway. Building on this insight, we address the TTA problem by leveraging high-frequency components in the data and propose a novel framework termed Frequency-based Decentralized Adaptation (FreDA). It partitions target samples into multiple homogeneous subdomains in the Fourier space, facilitating more accurate model adaptation. This is further enhanced by a frequency-based augmentation strategy that enriches each subdomain with diverse samples, strengthening model robustness.
V-A Frequency-based Decentralized Learning
Fourier transform offers an effective method to extract different frequency components from images, with high-frequency information particularly useful for capturing fine-grained details such as texture and noise. These details often highlight subtle variations among different distribution shifts. Based on this insight, we propose a module termed Frequency-based Decentralized Learning. It leverages frequency information extracted from the pixel space to systematically partition data into multiple homogeneous subsets, enabling multiple local models to specialize in capturing each distribution shift individually. Concurrently, our method enables collaborative learning by allowing periodic parameter sharing among these local models via an enhanced base model, thereby boosting overall adaptability to diverse distribution shifts.
@@ V-A4 Base Model Enhancement
To integrate knowledge from all subnetworks and prevent degradation on specific subdomains, we periodically aggregation their parameters at intervals of time $T$:
@@
where $|\mathcal{D}_{k}|$ denotes sample number in cluster $k$. This aggregation step combines the parameter updates from each subnetwork proportionally to its cluster size. The updated parameters $\mathbf{\theta}_{\text{base}}$ are then distributed back to each subnetwork, initializing them for the next batch of training: $\mathbf{\theta}_{k}\leftarrow\mathbf{\theta}_{\text{base}}$.

### v4

@@ Abstract
Deploying deep models in real-world scenarios remains challenging due to significant performance drops under distribution shifts between training and deployment environments. Test-Time Adaptation (TTA) has recently emerged as a promising solution, enabling on-the-fly model adaptation. However, its effectiveness deteriorates in the presence of mixed distribution shifts – common in practical settings – where multiple target domains coexist. In this paper, we study TTA under mixed distribution shifts and move beyond conventional whole-batch adaptation paradigms. By revisiting distribution shifts from a spectral perspective, we find that the heterogeneity across latent domains is often pronounced in Fourier space. In particular, high-frequency components encode domain-specific variations, which facilitates clearer separation of samples from different distributions. Motivated by this observation, we propose to un-mix heterogeneous data streams using high-frequency domain cues, making diverse shift patterns more tractable. To this end, we propose Frequency-based Decentralized Adaptation (FreDA), a novel framework that decomposes globally heterogeneous data stream into locally homogeneous clusters in the Fourier space. It leverages decentralized learning and augmentation strategies to robustly adapt under mixed domain shifts. Extensive experiments across various environments (corrupted, natural, and medical) show the superiority of our method over the state-of-the-arts.
@@ I Introduction
Despite their success, current TTA models are often limited to ideal testing conditions, typically involving homogeneous test samples with similar types of distribution shifts. In reality, distribution shifts are mixed, overlapping, and even conflicting [11, 12, 13, 14]. For instance, photo management software handles diverse corruptions, such as noise, blur, compression; medical imaging platforms manage inconsistencies from varied acquisition methods; and autonomous driving systems face fluctuating conditions like lighting, weather, and road types.
@@
Such complexity poses serious challenges for existing methods. While recent efforts have extended TTA to continually changing environments [5], they typically assume a uniform target domain at each time step, as illustrated in the continual setting in Fig. 1 (a). Some methods periodically reset the model to its source pre-trained state [7, 9] to mitigate the cumulative effect of sequential distribution shifts on adaptation. Others down-weight outlier target samples [6, 10] to suppress the impact of abrupt or rare shifts. Although these strategies aim to improve robustness under dynamic conditions, they lack the capacity to disentangle or localize domain variations within a batch. As a result, they often misalign conflicting domain shifts, failing to be deployed in realistic, heterogeneous data streams (see Fig. 1 (b)).
@@
Figure 1: (a) Illustration of different online distribution shifts, ranging from single-domain to continual and mixed-domain shifts, with increasing complexity and realism. (b) Classification error rate on CIFAR-10-C: while conventional models experience a sharp performance drop as distribution shifts become more complex, our method maintains superior performance, particularly under mixed-domain shifts.
@@
To address this, we propose shifting from coarse, homogeneous / whole-batch adaptation to a fine-grained strategy that explicitly disentangles heterogeneous shifts. Our approach leverages the spectral information to disentangle domain variations, which naturally separates sample features across frequency bands: high-frequency components capture fine details like edges and textures, while low-frequency components represent global structures such as shape and illumination (see Section IV). By decomposing inputs accordingly, we are able to better characterize distributional variations and identify diverse shifts. Moreover, the Fourier transform operates directly on raw pixel-level inputs, independent of any pre-trained model, making it robust to large domain gaps and enhancing adaptability in real-world deployments.
Figure 2: Illustration of FreDA framework. Heterogeneous data streams are partitioned in Fourier space into locally homogeneous clusters, where local models adapt independently and synchronize weights for intermittent update.
@@
Building upon this insight, we introduce Frequency-based Decentralized Adaptation (FreDA). FreDA begins by partitioning incoming data in the Fourier space, leveraging high-frequency components to group samples with similar domain characteristics. This transforms globally heterogeneous inputs into locally homogeneous subsets before any adaptation takes place. Multiple local models are then deployed, each assigned to a specific cluster, and adapt independently to their respective data streams while periodically uploading their parameters. These intermittent updates help build more stable model weights, reducing potential errors during clustering. To further enhance robustness, we introduce a Fourier-based augmentation scheme that improves sample quality and reinforces adaptation to shift-specific patterns. To summarize, the main contributions of this work are three-fold:
@@
  • We identify a key limitation of most existing TTA methods – their neglect of real-world data heterogeneity – which leads to suboptimal performance when confronted with diverse, mixed distribution shifts.
@@
  • We validate FreDA through extensive experiments on corrupted, natural, and medical benchmarks, demonstrating consistent improvements over state-of-the-art methods across various distribution shift scenarios.
@@ II-A Transfer Learning, Domain Adaptation, Test-time Training, and Test-time Adaptation
Deep neural networks often experience performance degradation when deployed in environments different from their training settings, due to distribution shifts between source (training) and target (testing) domains. This issue has been extensively studied under the umbrella of transfer learning [15, 16, 17], which aims to transfer knowledge across domains. A key branch of transfer learning is domain adaptation (DA), where models trained on the labeled source domain are adapted to the target domain. Depending on the data availability, previous DA can be roughly classified into: Supervised DA [18] using labeled data in both source and target domains during training; Unsupervised DA [19, 20, 21, 22], which relies on labeled source data and unlabeled target data simultaneously; and Source-free DA [23, 24, 25], that adapts to the target domain without access to source data, typically due to privacy constraints.
While traditional DA methods assume access to the entire target distribution before adaptation, real-world deployment often involves receiving data as an online stream, where samples arrive sequentially and the full distribution is never observed. This streaming nature of the target domain breaks the assumption of global access, making conventional domain adaptation infeasible. To handle this constraint, two prominent test-time adaptation paradigms have emerged. Online Test-time Training (OTTT) [26, 27] introduces an auxiliary task – often self-supervised – during pre-training. At test time, the model optimizes this auxiliary objective to adapt to the target distribution accordingly; Online Test-time Adaptation (OTTA) [3, 7, 6, 4, 5, 8] poses a more demanding and practical scenario, where the model must adapt on-the-fly to the test stream without any prior modification during training. This setting emphasizes the need for rapid, real-time model updates to effectively capture the continuously incoming data.
A brief comparison of these adaptation settings is summarized in TABLE I. Our work falls under the OTTA setting, where we propose a generalizable and robust adaptation framework that maintains strong performance across diverse distribution shifts and corruptions – without relying on source data or altering the pre-training process.
TABLE I: Comparison of different transfer learning settings where $\checkmark$ denotes settings closer to real-world conditions.
Research Fields No Source Data No Target Labels Online Adaptation Model Agnostic
Supervised Domain Adaptation $\times$ $\times$ $\times$
Unsupervised Domain Adaptation $\times$ $\checkmark$ $\times$
Source-free Domain Adaptation $\checkmark$ $\checkmark$ $\times$
Online Test-time Training $\checkmark$ $\checkmark$ $\checkmark$ $\times$
Online Test-time Adaptation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ II-B1 Mixed Domains
Modern TTA methods designed for continual domain shifts [5, 28, 6, 9, 10] have pushed beyond single adaptation problems. However, while they account for the dynamic target distribution shifts over time, they typically assume that each batch contains only a single domain shift. This simplifies the real-world scenario, where multiple distribution shifts may occur simultaneously, creating a mixed-domain environment that remains largely unexplored.
Existing approaches primarily focus on stabilizing model updates through periodic parameter resets [7, 9] or importance weighting of target samples [6, 9, 10], but these techniques are ill-equipped to disentangle and manage the complex interleaving of domain shifts, leading to degraded performance. In contrast, our work explicitly addresses mixed domains via frequency-space decomposition, which facilitates domain separation and proactive distribution alignment prior to model adaptation. While recent efforts like [7] acknowledge mixed distribution challenges under the broader “Dynamic Wild World” paradigm, their unified treatment of diverse real-world factors does not systematically address data heterogeneity. Conversely, our study directly targets the core issue of mixed domains in TTA, introducing tailored strategies for disentangling and adapting to co-occurring heterogeneous distributions.
@@ II-B2 Dependent Sampling
The second challenge arises from temporal-dependent sampling, which introduces class-level dependencies. This topic has garnered considerable research attention, with a growing body of work [28, 29, 30, 31, 32, 33] actively addressing it through strategies such as pseudo-label-based rebalancing or extended observation windows. These approaches primarily aim to address temporal imbalances caused by skewed class distributions over time, aiming to improve target domain estimation by addressing low class diversity within a batch. However, as these methods focus on within-domain temporal coherence, their applicability becomes limited when batches contain samples from multiple target domains. This limitation is also reflected in our experiments, where mixed-domain sampling poses notable challenges not accounted for in these approaches.
@@ III Problem Definition of mixed domain shifts
IV Un-mixing TTA under Mixed Distribution Shifts: A Fourier Perspective
While Test-Time Adaptation (TTA) methods perform well under single-type shifts, they degrade under mixed distribution shifts that are ubiquitous in real-world deployment (see Fig. 1 (b)). To address this, we advocate a paradigm of proactive distribution management: before any parameter update, we resolve heterogeneity by decomposing the input stream in frequency space. Our key insight is that different distributions leave distinct spectral footprints, which can be utilized to disentangle the hetergenous data stream. As shown in Fig. 3, high-frequency components naturally separate different domain shifts, in contrast to the substantial overlap observed in the latent sample features from a pre-trained model. This proactive step transforms an ill-posed, entangled adaptation problem into a set of well-conditioned homogeneous subproblems.
Take TBN [34] as an example. In its original centralized version, TBN computes statistical information using the whole target batch and directly replaces the source values for online adaptation. While effective under single-domain shifts, this whole-batch, homogeneous strategy produces biased statistics under mixed domain shifts, leading to suboptimal adaptation (see Fig. 4 (a)). In contrast, our decentralized approach first applies high-frequency clustering to partition the batch into locally homogeneous clusters. Batch normalization statistics are then computed within each cluster, enabling localized adaptation that can effectively prevent cross-domain contamination. Therefore, this decentralized strategy can reduce the error rate by a notable 4.27% and yields tighter within-class cohesion along with sharper between-class boundaries compared to centralized adaptation (see Fig. 4 (b)).
Figure 3: (a) Example histopathology patches from five distinct healthcare centers illustrate visual heterogeneity across different domains (Camelyon17 [35]). (b) Latent features from the pre-trained model fail to distinguish target subdomains and remain highly entangled. (c) In contrast, high-frequency features can clearly separate different target subdomains.
V Frequency-based Decentralized Adaptation
As outlined in the previous section, distinguishing samples from different distribution shifts provides a promising direction for TTA. Building on this insight, we introduce Frequency-based Decentralized Adaptation (FreDA), which partitions target samples in the Fourier space into homogeneous subdomains, enabling more accurate adaptation. We further enhance this process with a augmentation strategy that diversifies each subdomain in the Fourier space to improve model robustness.
V-A Frequency-based Decentralized Learning
We first introduce Frequency-based Decentralized Learning framework. It begins by extracting high-frequency information from the pixel space to partition target samples into homogeneous subsets, disclosing the latent target subdomains contained within a batch. We then perform a decentralized adaptation, where multiple local models – sharing the same architecture – independently adapt to their assigned subsets. At regular intervals, the parameters of these local models are aggregated to form a stronger base model, which is subsequently used to update the local models for the next adaptation phase. Detailed procedures are provided in the following subsections.
Methods CIFAR-10-C CIFAR-100-C ImageNet-C
@@ V-A Frequency-based Decentralized Learning
Figure 4: Classification error rate (%) and t-SNE sample feature visualization using TBN [34] under mixed distribution shifts. Different colors denote different classes. (a) Original Centralized Adaptation: global Batch Normalization parameters are computed from the whole target batch. (b) Our Decentralized Adaptation: local Batch Normalization parameters are tailored for clusters separated based on high-frequency features. Decentralized adaptation achieves a 4.27% error reduction and clearer class structure compared with centralized adaptation.
@@ V-A4 Base Model Enhancement
To mitigate potential clustering errors for more stable adaptation model weights, we allow each local models to periodically upload their parameters for aggregation at intervals of time $T$:
@@
where $|\mathcal{D}_{k}|$ denotes sample number in cluster $k$. This aggregation step combines the parameters from all local models in proportion to their cluster size. The updated parameters $\mathbf{\theta}_{\text{base}}$ are then distributed back to each subnetwork, initializing them for the next adaptation round: $\mathbf{\theta}_{k}\leftarrow\mathbf{\theta}_{\text{base}}$.